OpenStudy (anonymous):

Solve for x: 2 over 5 (x - 4) = 2x. -8 negative 1 over 2 1 -1

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

I think it is D @campbell_st

OpenStudy (campbell_st):

is the question \[\frac{2}{5}(x -4) = 2x\]

OpenStudy (anonymous):

Yes.

OpenStudy (campbell_st):

ok... so multiply both sides by 5 and you get 2(x - 4) = 10x so your solution is correct as 2x - 8 = 10x you can solve for x